OpenStudy (anonymous):

Does anyone know how to write this in sigma notation? -2, + 1 + 4 + ... + 13 …

OpenStudy (anonymous):

the answer my teacher gave me is \[\sum_{n=1}^{6}3n-5\] but can u help me understand how to get it?

OpenStudy (anonymous):

well when n=1 then 3-5=-2, n=2, then 6-5=1, n=3 then 9-5=4. when n=4 then 12-5=7, when n=5 then 115-5=10, when n=6 then 18-5=13

OpenStudy (anonymous):

did u get that?

OpenStudy (anonymous):

yes